OpenStudy (lilday12):

Convert 4.30 hours to minutes. Round answers to the nearest tenth. A.258 minutes B.14.0 minutes C.15,480 minutes

OpenStudy (abdullahm):

1 hour = 60 minutes 4.3 hours = ?? minutes Multiply both sides by 4.3 1 hour * 4.3 = 60 * 4.3 minutes 4.3 hours = ?? minutes

OpenStudy (abdullahm):

So basically, what is 4.3 * 60 = ?